# BlockingPy: approximate nearest neighbours for blocking of records for entity resolution

### v1

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which utilises blocking via modern approximate nearest neighbour search and graph algorithms to significantly reduce the number of comparisons. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. We believe that the presented software will be useful for researchers (i.e., social scientists, economists or statisticians) interested in linking data from various sources.
@@
C1 Current code version v0.1.14
@@
C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack
@@ 1 Motivation and significance
Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers. This process underpins countless research endeavours across disciplines including epidemiology, economics, official statistics or historical research where data integration is fundamental to scientific discovery [1, 2]. Traditional blocking techniques, while computationally efficient, exhibit significant limitations when confronted with real-world data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting. These challenges, thoroughly documented by [3, 4, 5], have not been adequately addressed by existing software solutions, which often force researchers to choose between computational feasibility and linkage accuracy.
The scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], or fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12]. However, these solutions typically rely on user-defined deterministic rules that may miss certain group of records or result in a large blocks of records when data irregularities are present (with the exception of the klsh, an R package that relies on the Locality Sensitive Hashing; [13]).
The BlockingPy package distinguishes itself by using state-of-the-art approximate nearest neighbour (ANN) search algorithms and graph-based indexing techniques that maintain robustness even with the imperfect data. This combination of modern algorithmic approaches represents a significant advancement over the fragmented workflow imposed by current tools, providing researchers with a seamless end-to-end pipeline that adaptively manages blocking strategies based on data characteristics. Users interact with BlockingPy through an user-friendly Python API that integrates with standard packages (e.g. pandas; cf. [14, 15]), allowing researchers to focus on analytical questions rather than technical implementation details of record linkage.
The structure of the paper is as follows. In Section 2 we briefly review the blocking procedures in existing Python packages. In Section 3 we provide details regarding the implementation ant the API. In Section 4 we provide two examples for probabilistic record linkage and deduplication. The paper finishes with brief review of the possible impact and conclusions.
@@ 2 Existing software covering blocking procedures
In this section, the focus is on Python, but readers are referred to the fastLink or reclin2 packages for more information on deterministic blocking procedures.
The splink package allows for deterministic blocking rules using function block_on() which specifies either the variables (e.g. block_on("city")) or more advanced combinations (e.g. block_on("substr(first_name, 1,1)", "surname")) which then are translated to SQL query for the DuckDB (via the BlockingRuleCreator class). More advanced users may be interested in creating their own rules which is possible through the CustomRule class. The splink allows users to specify the list of blocking rules to the Linker class. The splink allows users to inspect the blocking via the
@@
The recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods. This package allows for deterministic blocking by BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by vector of column names and optionally window for the SortedNeighbourhood() which accounts for multiple mistakes.
The only software that provides similar functionalities is the blocklib package, which is part of the Anonlink project for privacy preserving record linkage (cf. [16, 17]). Unfortunately, at the time of writing, it has been two years since the last update, and this package no longer appears to be supported. We should also mention two recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, and semhash [19] that allows deduplication of records using the ANN algorithms via the vicinity. However, none of these packages provides an easy way to use this information for blocking of records for probabilistic record linkage and deduplication. Thus the BlockingPy provides a significant extensions to the existing Python packages.
@@ 3.1 Installation
BlockingPy is a Python package released under the MIT license. It can be installed from PyPI with popular package managers, e.g., pip or Poetry.
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input.
Subsequently, similarity-based nearest neighbour search is being performed for each record in either the input dataset (deduplication) or query dataset (record linkage). This step is accomplished using one of the following Python implementations of ANN algorithms:
@@
Users can fine-tune each algorithm according to their needs. Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package; [26]), generated by previous neighbourhood search results. Figure 1 illustrates this workflow. By using the block() method with desired parameters on the Blocker instance, users can incorporate the above-mentioned process to obtain blocking results.
Figure 1: Blocking workflow in the BlockingPy package
@@ 3.2.2 Evaluation of blocking result
The blocking quality can be evaluated by providing ground truth blocks to either the block() or eval() methods. This allows users to assess the quality of the blocks computed by BlockingPy. The package implements batch processing for evaluation to allow processing of memory-intensive evaluations. The key evaluation information generated by the package is described in the sections below.
@@
where $k$ is the total number of blocks, $n$ is the total number of records in the dataset, and $|B_{i}|$ is the number of records in the $i$-th block. $\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}$ is the number of comparisons after blocking, while $\binom{n}{2}$ is the total possible comparisons without blocking. For record linkage reduction rate is defined as follows
@@
Confusion Matrix. Presents results in comparison to ground-truth blocks in a pairwise manner (e.g., one true positive pair occurs when both records from the comparison pair belong to the same predicted block and to the same ground-truth block in the evaluation data frame.).
@@
Evaluation metrics. Enables users to evaluate blocking across metrics as follows:
  • 1.
    Recall – $\frac{TP}{TP+FN}$.
  • 2.
    Precision – $\frac{TP}{TP+FP}$.
  • 3.
    Accuracy – $\frac{TP+TN}{TP+TN+FP+FN}$.
  • 4.
    Specificity – $\frac{TN}{TN+FP}$.
  • 5.
    F1 Score – $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$.
  • 6.
    False Positive Rate – $\frac{FP}{FP+TN}$.
  • 7.
    False Negative Rate – $\frac{FN}{FN+TP}$.
@@ 3.3 Software architecture
The architecture of BlockingPy provides a modular design, where each component is responsible for a specific aspect of the entire blocking workflow. This approach facilitates the integration of new ANN algorithms and modifications to existing functionality. The main components are illustrated in Figure 2 and their roles are described below.
@@
Blocker. The primary component of the BlockingPy architecture. It coordinates interaction between other components, managing the whole workflow of the package from handling data ingestion to generating results.
ANN Algorithms. Collection of ANN implementations with common interface for the straightforward addition of new algorithms. The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method. An example of this is shown in Listing 1.
@@
Input Processing. Manages the data pre-processing and transformation across multiple input formats. Converts dense and sparse matrices or raw text data to sparse data frames for efficient processing. Currently, we support transformation based on $n$-grams but in the future developments we plan to include support for vectors as in the semhash package.
Controls. Manages the configuration system for both text and ANN parameters. Includes customization for algorithm-specific options e.g., metric selection, search and query parameters, alongside tokenization rules and $n$-gram size. Listing 2 presents an example of Controls dictionary for the hnsw algorithm.
@@
Input Validation. Ensures data and parameter correctness, while providing error messages for the user.
Blocking Result. Component responsible for management of the results and the corresponding information about the blocking process e.g., algorithm used, evaluation metrics, reduction ratio, or block distribution.
@@ 4 Illustrative examples
In the following section, we demonstrate BlockingPy package’s functionality on two examples of both record linkage and deduplication tasks. We aim to show users how user-friendly and effective the package is when blocking records is involved in entity resolution pipelines.
@@ 4.1 Record linkage
In this example, we use the Cis-Census datasets [27] to show the record linkage capabilities of BlockingPy. These datasets contain fictitious personal information, with the census dataset comprising 25,343 records and the cis (Customer Information System) dataset containing 24,613 records. Table 2 shows sample records from the census dataset after initial preparation conducted in Listing 3 for reference and the cis dataset is structured similarly.
@@
Subsequently, in Listing 5 we initialise Blocker and use the block() method and select the hnsw algorithm with default parameters to obtain the results. Tokenization and white-space removal are handled by the algorithm internally and can be modified via the control_txt parameter and the algorithm parameters can be fine-tuned via the control_ann argument.
@@
The output is a data frame containing rows with indices from both datasets (x, y) alongside the assigned block (an integer starting with 0) and distance computed with one of the available metrics. Listing 6 presents a snippet of the output in this example and Listing 7 shows the basic blocking information logged by the package where the first section informs about the algorithm used, the number of generated blocks, the number of features (created by shingling), the reduction ratio, and the second section displays the counts of blocks with each size, e.g., there were $591$ blocks created with each containing $3$ records.
@@
5Number of columns used for blocking: 1072
@@
We observe a reduction ratio of $0.9999$, which indicates that the possible number of candidate pairs is substantially reduced, thus lowering the computational complexity which is crucial for large-scale datasets.
With a unique personal identifier available in both datasets, we can evaluate the algorithm. For that, we create true_blocks data frame which represents indices from both datasets and their true block, as shown in Listing 8. In this case, every record in each block should point to the same unique entity. Furthermore, in Listing 9 we put the rec_lin_result object (from Listing 5) and the true_blocks in the eval method to obtain the evaluation results. In this example, we randomly sampled 1000 ground-truth pairs to showcase the evaluation functionality of the BlockingPy package.
@@
Through this step, the user can access both the metrics and the confusion matrix. Evaluation is performed only for records that are available in true_blocks. We assume that we do not have knowledge about the records not included in the ground-truth data frame. Listing 10 and Listing 11 present the metrics and confusion matrix computed in this example.
@@ 4.2 Deduplication
In this example, our aim is to show how BlockingPy can be integrated into full entity resolution workflows. For that, we will present the deduplication of febrl1 dataset obtained from the recordlinkage package [10]. This dataset was generated using the FEBRL software [3] and contains 1000 records with fictitious personal information, of which 500 are original and 500 are duplicates with introduced errors. Firstly, we prepare the blocking key by imputing missing values with an empty string and concatenating fields with personal information into a single field, which we will use for blocking. After initialising Blocker and performing blocking on said data we are given the blocks by BlockingPy. Listing 12 presents this process, and the Listing 13 shows the outputted blocking information and the first five rows of the actual result.
@@
Both columns x and y refer to indices from the febrl1 alongside their designated block and distance measured between them with one of the available metrics (cosine in this example, the default for hnsw). Furthermore, the necessary step to make integration with recordlinkage possible is to add the block column to the original data frame, which we can do by transforming the blocking result as presented in Listing 14.
1result_df = blocking_result.result
2
3mapping_df = (
4 result_df
5 .melt(id_vars=[’block’], value_vars=[’x’, ’y’], value_name=’record_id’)
6 .drop_duplicates(subset=[’record_id’])
7)
8record_to_block = dict(zip(mapping_df[’record_id’], mapping_df[’block’]))
9new_data = df.copy()
10new_data[’block’] = [record_to_block.get(i) for i in range(len(df))]
@@
The full version of this example alongside the code can be found in the documentation.
@@
3pairs = indexer.index(new_data)
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one data set. For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of the residents of Ukraine under temporary protection in Poland as of March 31, 2023 [28]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr).
In our opinion, this software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist. Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools. The software has significantly improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records. Prior to this solution, Statistics Poland’s staff struggled with linking data without identifiers, applying user-defined rules which may have led to potential overestimation or underestimation of population flows.
Therefore, we believe that the proposed software will be of interest for researchers and practitioners who link data without identifiers or would like to apply large language models to detect duplicates but need to reduce costs by providing a small number of comparison buckets.

### v2

@@ Abstract
Entity resolution (probabilistic record linkage, deduplication) is a key step in scientific analysis and data science pipelines involving multiple data sources. The objective of entity resolution is to link records without common unique identifiers that refer to the same entity (e.g., person, company). However, without identifiers, researchers need to specify which records to compare in order to calculate matching probability and reduce computational complexity. One solution is to deterministically block records based on some common variables, such as names, dates of birth or sex or use phonetic algorithms. However, this approach assumes that these variables are free of errors and completely observed, which is often not the case. To address this challenge, we have developed a Python package, BlockingPy, which uses blocking via modern approximate nearest neighbour search and graph algorithms to reduce the number of comparisons. In this paper, we present the design of the package, its functionalities and two case studies related to official statistics. The presented software will be useful for researchers interested in linking data from various sources.
@@
C1 Current code version v0.1.15
@@
C7 Compilation requirements, operating environments & dependencies pandas, numpy, scipy, annoy, hnswlib, pynndescent, scikit-learn, networkx, nltk, voyager, faiss-cpu, mlpack, model2vec
@@ 1 Motivation and significance
Entity resolution requires statistical and computational methodologies to accurately identify matching records across datasets without unique identifiers. This process underpins countless research endeavours in disciplines such as epidemiology, economics, official statistics and historical research where data integration is fundamental to scientific discovery [1, 2]. Traditional blocking techniques, while computationally efficient, exhibit limitations when confronted with various data quality issues such as missing values, transliteration, typographical errors, and inconsistent formatting. Existing deterministic blocking methods rely on user-defined exact-matching rules that assume key fields (e.g., name, date of birth) are error-free, potentially leading to missed links and false matches [3, 4, 5].
Scientific community has developed several tools to address record linkage, including R packages such as RecordLinkage [6], reclin2 [7], fastLink [8], or Python implementations like splink [9], recordlinkage [10], Dedupe [11] or FEBRL [12]. However, these solutions typically rely on user-defined deterministic rules that may miss certain groups of records or result in large blocks of records when data irregularities are present.
Some solutions regarding these issues were proposed in the literature, for instance, the use of the SOUNDEX algorithm and other phonetic algorithms to block records on character strings that sound similar phonetically [13]; blocking using the Locality Sensitive Hashing (LSH; [14]), an approximate nearest neighbour (ANN) algorithm to search for similar records (see klsh package in R) or usage of Bloom filters, useful in particular for privacy-preserving record linkage (see, [15]; for implementation see blocklib in Python).
The BlockingPy package distinguishes itself by using state-of-the-art ANN search algorithms and graph-based indexing techniques, which allow it to work even with imperfect data. BlockingPy implements an ANN-based blocking procedure ($n$-gram encoding or embeddings, neighbour search, graph components) behind a single pandas-style call, cutting custom rule-writing and setup time.
The structure of the paper is as follows. In Section 2, we briefly review the blocking procedures in existing Python packages. In Section 3, we provide details regarding the implementation and the API. In Section 4, we provide two examples of probabilistic record linkage and deduplication. The paper finishes with a brief review of the possible impact and conclusions.
@@ 2 Existing software covering blocking procedures
In this section, the focus is on Python, but readers are referred to the fastLink and reclin2 packages for more information on deterministic blocking procedures.
The splink package allows for deterministic blocking rules using the function block_on() which specifies either the variables (e.g. block_on("city")) or more advanced combinations (e.g. block_on("substr(first_name, 1,1)", "surname")), which then are translated to a SQL query for the DuckDB (via the BlockingRuleCreator class). More advanced users may be interested in creating their own rules, which is possible through the CustomRule class. Splink allows users to specify the list of blocking rules for the Linker class. Splink allows users to inspect the blocking via the
@@
Additionally, splink supports blocking using phonetic algorithms such as Soundex, Metaphone, or Double Metaphone either by applying them directly within the blocking rules (e.g., block_on("soundex(surname)")), or by preprocessing columns to store phonetic encodings separately.
Recordlinkage package supports deterministic blocking via the recordlinkage.Index class and its methods. This package allows for deterministic blocking using BlockIndex() and SortedNeighbourhood() where specification of blocking variables should be provided by a vector of column names and optionally window for the SortedNeighbourhood(), which accounts for multiple mistakes.
The only software that provides similar functionality is the blocklib package, which is part of the Anonlink project for privacy-preserving record linkage (cf. [16, 17]). We should also mention three recent packages: vicinity [18], which provides a similar API for the ANN algorithms described in Section 3, semhash [19], which allows deduplication of records using the ANN algorithms via the vicinity and LinkTransformer [20], which uses kNN-based embedding comparisons and offers exact-match blocking. However, none of these packages provide a way to use this information for blocking of records for probabilistic record linkage and deduplication. Thus, BlockingPy provides an extension to the existing Python package ecosystem.
@@ 3.1 Installation
BlockingPy is a Python package released under the MIT license. It can be installed from PyPI using package managers such as pip or Poetry.
@@ 3.2 Software functionalities
This subsection describes the main functionalities provided by the BlockingPy package.
@@ 3.2.1 Blocking records
The core functionality of BlockingPy is to provide an efficient and scalable approach for blocking (also known as indexing) records in both record linkage and deduplication pipelines. To achieve this, users can input either previously computed Document-Term Matrices (DTMs) or raw text data, which the package transforms into DTMs by constructing character $n$-gram representations of the input. Alternatively, users can choose to represent the input data as vector embeddings.
Subsequently, similarity-based nearest neighbour search is performed for each record in either the input dataset (deduplication) or query dataset (record linkage). This step is accomplished using one of the following Python implementations of ANN algorithms:
@@
Users can fine-tune each algorithm according to their needs. Finally, BlockingPy groups records into blocks by identifying connected components in an undirected graph (using the igraph package [27]), generated by previous neighbourhood search results. By using the block() method with the desired parameters on the Blocker instance, users can run the above-mentioned process to obtain blocking results.
@@ 3.2.2 Evaluation of blocking result
The blocking quality can be evaluated by providing ground truth blocks to either the block() or eval() methods. This allows users to assess the quality of the blocks computed by BlockingPy. The key evaluation information generated by the package is described in the sections below.
@@
where $k$ is the total number of blocks, $n$ is the total number of records in the dataset, and $|B_{i}|$ is the number of records in the $i$-th block. $\sum\limits_{i=1}^{k}\binom{|B_{i}|}{2}$ is the number of comparisons after blocking, while $\binom{n}{2}$ is the total number of possible comparisons without blocking. For record linkage the reduction ratio is defined as follows
@@
Confusion matrix presents results in comparison to ground-truth blocks in a pairwise manner (e.g., one true positive pair occurs when both records from the comparison pair belong to the same predicted block and to the same ground-truth block in the evaluation data frame).
@@
Finally, table 2 presents metrics that enables users to evaluate blocking procedures.
Metric Formula Metric Formula
Recall $\frac{TP}{TP+FN}$ Accuracy $\frac{TP+TN}{TP+TN+FP+FN}$
Precision $\frac{TP}{TP+FP}$ Specificity $\frac{TN}{TN+FP}$
F1 Score $2\cdot\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ False Positive Rate $\frac{FP}{FP+TN}$
False Negative Rate $\frac{FN}{FN+TP}$
Table 2: Evaluation Metrics
@@ 3.3 Software architecture
The architecture of BlockingPy provides a modular design, where each component is responsible for a specific aspect of the entire blocking workflow. This approach facilitates the integration of new ANN algorithms and modifications to existing functionality. The main components are illustrated in Figure 1, and their roles are described below.
@@
Blocker. The primary component of the BlockingPy architecture. It coordinates interactions between other components, managing the whole workflow of the package from handling data ingestion to generating results.
ANN Algorithms. Collection of ANN implementations with a common interface for the straightforward addition of new algorithms. The parameters of each algorithm can be adjusted by providing the Controls dictionary to the control_ann parameter within the block method. An example of this is shown in Listing 1.
@@
Input Processing. Manages the data pre-processing and transformation across multiple input formats. Converts dense matrices, sparse matrices, and raw text data to sparse data frames for efficient processing. Currently, we support transformation based on $n$-grams and vector embeddings.
Controls. Manages the configuration system for both text and ANN parameters. Includes customization for algorithm-specific options e.g., metric selection, search and query parameters, alongside tokenization rules and $n$-gram size, or alternatively the choice of an embedding model. Listing 2 presents an example of a Controls dictionary for the hnsw algorithm.
@@
Input Validation. Ensures validity of the data and parameters while providing error messages for the user.
Blocking Result. The component responsible for managing results and the corresponding information about the blocking process e.g., algorithm used, evaluation metrics, reduction ratio, and block distribution.
@@ 4 Illustrative examples
In the following section, we demonstrate the BlockingPy package’s functionality using two examples: one for record linkage and one for deduplication. We aim to show the ease of use and effectiveness of the package in blocking records for entity resolution.
@@ 4.1 Record linkage
In this example, we use the Cis-Census datasets (for more information see [28]) to show the record linkage capabilities of BlockingPy. These datasets contain fictitious personal information, with the census dataset comprising 25,343 records and the cis (Customer Information System) dataset containing 24,613 records. Table 3 shows sample records from the census dataset after the initial preparation described in Listing 3. The cis dataset is structured similarly.
@@
Subsequently, in Listing 5 we initialise Blocker, invoke the block() method, and select the hnsw algorithm with default parameters to obtain the results. The tokenization and white-space removal are handled by the algorithm internally and can be modified via the control_txt parameter. Additional algorithm-specific parameters can be fine-tuned via the control_ann argument.
@@
The output is a data frame containing rows with indices from both datasets (x, y) alongside the assigned block (an integer starting with 0) and distance computed with one of the available metrics. Listing 6 presents a snippet of the output in this example. Listing 7 shows the basic blocking information logged by the package. The first section provides details about the algorithm used, the number of generated blocks, the number of columns used for blocking (created by shingling, a preprocessing step that converts strings into overlapping letter groups used as input to the ANN algorithm), and the reduction ratio. The second section displays the distribution of block sizes; for example, there were $591$ blocks created, each containing $3$ records.
@@
5Number of columns created for blocking: 1072
@@
We observe a reduction ratio of $0.9999$, which indicates that the possible number of candidate pairs is substantially reduced, thus lowering the computational complexity, which is crucial for large-scale datasets.
With a unique personal identifier available in both datasets, we can evaluate the algorithm. For that, we create a true_blocks data frame which represents indices from both datasets and their true block, as shown in Listing 8. In this case, every record in each block should point to the same unique entity. Furthermore, in Listing 9 we pass the rec_lin_result object (from Listing 5) and the true_blocks to the eval method to obtain the evaluation results. In this example, we randomly sampled 1000 ground-truth pairs to showcase the evaluation functionality of the BlockingPy package.
@@
Through this step, the user can access both the evaluation metrics and the confusion matrix. Evaluation is performed only for records that are available in true_blocks. We assume that we have no knowledge about the records not included in the ground-truth data frame. Listing 10 and Listing 11 present the metrics and confusion matrix computed in this example.
@@ 4.2 Deduplication
In this example, our aim is to show how BlockingPy can be integrated into full entity resolution workflows. For that, we will present the deduplication of the febrl1 dataset obtained from the recordlinkage package [10]. This dataset was generated using the FEBRL software and contains 1000 records with fictitious personal information, of which 500 are original and 500 are duplicates with introduced errors. Firstly, we prepare the blocking key by imputing missing values with an empty string and concatenating fields with personal information into a single field, which we will use for blocking. After initialising Blocker and performing blocking on said data, we are given the blocks by BlockingPy. Listing 12 presents this process, and the Listing 13 shows the blocking information output and the first five rows of the actual result.
@@
Both columns x and y refer to indices from the febrl1 dataset alongside their designated block and distance measured between them using one of the available metrics (cosine in this example, the default for hnsw). Furthermore, the necessary step to make integration with recordlinkage possible is to add the block column to the original data frame, which we can do with the add_block_column method as presented in Listing 14.
1df[’id’] = range(len(df))
2df_final = blocking_result.add_block_column(df, id_col_left=’id’)
@@
The full version of this example and many others, alongside the code, can be found in the documentation.
@@
3pairs = indexer.index(df_final)
@@ 5 Impact
The proposed approach is used at Statistics Poland in the procedure for the linkage of records between administrative datasets as well as within one dataset. For instance, we have used this approach to deduplicate the number of forced migrants crossing the Polish-Ukrainian border after the full-scale invasion of Ukraine by Russia on February 24, 2022. We also used this technique to determine the number of residents of Ukraine under temporary protection in Poland as of March 31, 2023 [29]. The main motivation behind this was the problem of lack of identifiers as well as the transliteration from Ukrainian/Russian to Polish language (e.g. Олександр, should be Ołeksandr not Aleksandr).
This software opens up new research questions regarding optimal blocking strategies in administrative data integration, particularly in contexts where language differences and transliteration complexities exist. Researchers can now investigate how different blocking approaches affect linkage quality in multilingual settings, a question that was previously difficult to study systematically without appropriate tools. The software has meaningfully improved the pursuit of quality of official statistics by allowing for more accurate population estimates through better deduplication of records. Before implementing this solution, Statistics Poland relied on user-defined rules for linking data without identifiers, which often resulted in imprecise population flow estimates due to missed matches or false positives.
We believe that the proposed software will be of interest to researchers and practitioners who link data without identifiers or would like to apply large language models to detect duplicates but need to reduce costs by providing a small number of comparison buckets.